# The Coherence Trap: When MLLM-Crafted Narratives Exploit Manipulated Visual Contexts

### v1

@@ 3.4 Dataset Statistics
With the above steps, we finally harvest our MDSM dataset $\mathcal{D}_{M}$, a large scale, semantic-aligned multi-modal benchmark with high-fidelity texts from MLLM. The detailed distribution of MDSM data is shown in Fig. 2(a). Compared with the existing manipulation detection benchmarks in Tab. 1, MDSM has the following advantages: 1)Risk Consideration of MLLM. MDSM acknowledges the emerging challenges posed by MLLMs and utilizes multi-modal to create semantically coherent and contextually plausible narratives for manipulated images. This scenario, though under-explored, is highly significant for the detection and grounding of manipulated multi-modal media. 2)Semantic Alignment. MDSM is an aligned image-text multi-modal media manipulation benchmark, which is a significant and more practical scenario for multi-modal manipulation detection. 3)Large Scale. Our MDSM comprises 441,423 samples and is the largest benchmark for detecting and grounding multi-modal manipulation. 4)Diverse Multi-media Sources. The multi-modal media of MDSM sources from diverse media sites, including The Guardian, The New York Times, The Washington Post, USA Today, and the BBC. Consequently, the generality of methods can be assessed via cross-domain evaluation.
@@ 4 Methodology
Figure 3: Illustration of our proposed AMD. Given a manipulated image-text pair, (a) maps the image and prompts into a unified dimension and concatenates them with the artifact token, (b) extracts perceptual clues into the artifact token, (c) integrates multi-modal features to generate text-based detection results, and (d) outputs and visualization.
Fig. 3 illustrates the architecture of our Artifact-aware Manipulation Diagnosis framework (AMD). Built upon Florence-2 [4] to leverage real-world knowledge, AMD follows a sequence-to-sequence framework for joint textual detection and grounding. Multimodal inputs are processed through three stages, as shown in Fig. 3a-c, and outputs localized predictions with textual explanations.
@@
In this prompt paradigm,
[…]
  • $[<\textit{Coordinates}>]$: Optionally encloses tampered region coordinates if the image is altered.
Artifact Token Embeddings. To effectively adapt the MLLM into MDSM context while preserving its pretrained knowledge, we introduce a learnable Artifact Token that explicitly encodes artifacts from heterogeneous inputs. Formally, let the artifact token embeddings be denoted as ${E}_{a}\in\mathbb{R}^{n_{a}\times d}$, where $n_{a}$ indicates the token count and $d$ the embedding dimension. The textual input is processed through an embedding layer to obtain text embeddings ${E}_{t}\in\mathbb{R}^{n_{t}\times d}$ , while the visual input is encoded via a vision backbone $\mathcal{E}_{v}$ followed by a LayerNorm-augmented linear projection, yielding image embeddings ${E}_{v}\in\mathbb{R}^{n_{v}\times d}$. The above embeddings are concatenated to construct the input sequence: ${S}_{\text{inp}}=[{E}_{v};{E}_{a};{E}_{t}]$, where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] means concatenating along the token dimension.
@@ 4.2 Artifact Pre-perception Encoding
This stage aims to perceive manipulation artifacts within input data and condense these forensic clues into the artifact token. Specifically, the input sequence $S$ undergoes processing through the multimodal encoder $\mathcal{E}^{p}_{m}$, yielding $\hat{S}=[\hat{E}_{v};\hat{E}_{a};\hat{E}_{t}]$. To inject artifact-aware clues into the artifact token embedding $\hat{E}_{a}$, we pick $\hat{E}_{a}$ from $\hat{S}$ and feed it into an artifact-aware classification head. As illustrated in Fig. 3, this classification head is optimized via a manipulation detection objective to explicitly encode artifact-related patterns into $\hat{E}_{a}$.
@@
$\mathcal{L}_{APE}=\mathbb{E}_{(I,T)\sim\mathcal{D}_{M}}\mathbf{CE}(C_{a}(\bar{E}_{a}),y_{\text{fd}}),$ (2)
@@ 4.3 Manipulation-Oriented Reasoning
Manipulation-Oriented Reasoning (MOR) is in charge of generating the textual answer in response to the question prompt. MOR first encodes the sequence ${S}_{a}$, subsequently, the language $Decoder$ takes the resultant features to generate the answers. To acquire an accurate response, we augment the network optimization in MOR with two guiding tasks: visual Artifact Capture via Grounding and Manipulation-focused Guidance.
@@
Subsequently, the $u_{agg}$ is sent to the bbox detector to generate artifact coordinates. We follow [34] to construct the image manipulation grounding loss using L1 loss $\mathcal{L}_{1}$ and GIoU loss $\mathcal{L}_{IoU}$:
@@
Language modeling. The input sequence $S_{a}$ is processed through an encoder-decoder architecture, ultimately generating a pure text output that includes choices and coordinates (Fig. 3d) as specified in the prompts. In this stage, an autoregressive approach is adopted, where the decoder generates the target sequence $y$ conditioned on $S_{m}$. The language modeling loss $\mathcal{L}_{LM}$ [4] is used to supervise the decoded text outputs.
@@ 4.5 Training and Inference
5 Experiments
Table 2: Comparison of multi-modal learning methods for MDSM, where the background gray indicates the intra-domain performance. The better results in each group are in boldface. AVG refers to the average performance across the five news domains.
@@ 5 Experiments
Zero-Shot Qwen2.5-VL [3] 47.74 29.24 0.00 35.18 25.70 0.00 24.66 40.60 0.00 25.11 40.29 0.28 35.89 31.51 0.00 33.72 33.47 0.06
Qwen3-a22b [43] 45.29 25.01 0.71 38.12 27.41 1.19 22.87 39.17 0.10 22.87 40.77 1.34 37.22 31.60 0.98 33.27 33.69 0.86
GPT-4o [14] 48.48 27.90 0.82 35.68 29.49 1.23 24.62 39.88 1.37 23.62 38.89 1.22 37.19 30.48 1.20 33.92 33.33 1.17
Gemini-2.0 [42] 56.05 33.16 1.44 41.26 24.37 1.12 29.60 38.29 1.40 29.15 35.20 2.42 38.12 29.13 2.25 38.83 32.03 1.72
Tr. on NYT ViLT [16] 83.27 64.27 22.73 72.18 31.76 20.21 70.34 36.45 21.48 65.71 36.23 17.56 74.33 36.10 19.36 73.17 40.96 20.27
HAMMER [38] 79.20 55.86 51.34 68.23 40.10 21.56 71.52 41.17 13.74 68.50 41.47 13.92 67.37 42.23 16.12 70.96 44.16 23.34
HAMMER++ [39] 79.61 57.06 54.44 66.99 38.07 17.34 67.18 37.58 10.76 66.28 37.97 10.88 66.12 37.82 13.68 69.23 41.70 21.42
FKA-Owl [22] 94.67 78.18 55.81 77.20 46.88 43.67 78.00 44.45 50.73 75.49 50.83 43.53 84.65 60.73 43.28 81.60 56.77 46.23
@@
Tr. on Guardian ViLT [16] 68.80 43.99 21.77 85.29 67.34 41.80 70.34 46.24 37.68 78.61 47.17 38.13 80.00 44.79 38.97 76.61 49.90 35.67
HAMMER [38] 61.89 37.98 18.84 78.50 52.40 51.53 74.78 50.76 43.40 75.11 50.34 46.36 81.32 50.15 56.03 74.32 48.33 43.23
HAMMER++ [39] 62.75 36.45 23.76 80.95 59.92 64.67 75.36 48.77 47.13 76.30 49.56 48.91 80.12 50.36 57.97 75.10 49.01 48.49
FKA-Owl [22] 80.60 40.44 26.33 92.60 78.24 71.04 80.90 51.80 50.93 78.88 51.62 50.88 87.61 68.57 61.82 84.12 58.13 52.20
@@
Table 3: Comparison of multi-modal learning methods for $\text{DGM}^{4}$, where the guardian domain with background gray is intra-domain. $\text{P}_{tok}$ is Precision of fake token grounding.
@@
ACC mAP $\text{P}_{tok}$ mIoU ACC mIoU $\text{P}_{tok}$ mIoU ACC mAP $\text{P}_{tok}$ mIoU ACC mAP $\text{P}_{tok}$ mIoU ACC mAP $\text{P}_{tok}$ mIoU
ViLT [16] 68.27 42.29 69.87 43.19 52.79 31.28 62.11 33.78 55.76 33.26 57.17 31.10 44.14 39.68 59.06 21.96 55.24 36.63 62.05 32.49
HAMMER [38] 78.34 66.79 78.27 61.09 64.97 40.49 73.76 40.51 63.54 40.26 76.13 38.53 54.97 40.84 81.48 43.74 65.45 47.10 77.41 45.97
HAMMER++ [39] 79.13 67.11 78.24 62.15 65.25 40.74 73.24 41.14 63.83 40.34 76.17 38.21 54.24 41.25 81.73 43.23 65.61 47.36 77.34 46.19
FKA-Owl [22] 82.97 53.86 87.70 65.69 67.57 38.97 79.44 32.57 67.05 37.70 81.55 31.86 70.26 40.20 84.54 46.48 71.96 42.68 83.31 44.15
@@
Experimental Setup. DaViT [9] serves as the image encoder $\mathcal{E}_{v}$. The multimodal encoder $\mathcal{E}_{m}$ and decoder are based on Florence-2[4], and $\mathcal{E}_{m}^{p}$ is kept frozen during training. The length of Artifact Token is set to 32. Code will be released for reproducibility.
5.1 Quantitative Results
Effectiveness & Generalization. We assess AMD’s in-domain and cross-domain performance against four SOTA methods on the MDSM and $\text{DGM}^{4}$ dataset. For MDSM, models are trained on two larger subsets, The Guardian and The New York Times (NYT), and tested on the remaining ones (Tab. 2). For $\text{DGM}^{4}$, we use the largest subset, The Guardian, as the training domain (Tab. 3). Additionally, Tab. 2 presents zero-shot results of leading general-purpose models on MDSM. Evaluation metrics are following [38]. Our key observations are as follows:
1) Forgery-trace-based methods, ViLT [16] and HAMMER [38], show limited performance within and across domains. In contrast, MLLM-based approaches, including FKA-Owl [22] and AMD, achieve significantly stronger results. For example, trained on MDSM-NYT (Tab. 2), AMD attains an intra-domain mAP of 84.47 and cross-domain scores above 60, while HAMMER has an mAP of 55.86 and cross-domain performance below 45. This affirms that MLLMs’ open-world knowledge enhances manipulation detection in complex scenarios.
Table 4: Ablation on each components (a) and discussion regarding performance on test set of difference MLLMs & differnece manipulation type (b).
@@ 5.1 Quantitative Results
2) On both MDSM and $\text{DGM}^{4}$, AMD consistently achieves strong grounding performance. In tab. 2, when trained on NYT and Guardian, AMD achieves average mIoU of 63.56 and 61.02, respectively, while FKA-Owl attains only 46.23 and 52.20. Although we designed more detailed prompts for general-purpose models, their grounding performance remains extremely poor, with mIoU scores not exceeding 3. This superiority of AMD stems from our carefully crafted question-answer heuristic prompts and the unified decoding framework of the MOR module. Unlike other methods that produce grounding results for every sample, AMD omits coordinate outputs when no manipulation is detected, thereby avoiding unnecessary errors and alleviating performance bottlenecks.
3) On $\text{DGM}^{4}$ (Tab. 3), MLLM-based approaches also demonstrate superior performance. AMD achieves 74.47 ACC, 52.91 mAP, 80.01 $\text{P}_{tok}$, and 51.87 mIoU, outperforming the HAMMER series. Despite trailing FKA-Owl in new text localization task metric $\text{P}_{tok}$, AMD surpasses it in the other three tasks, especially mAP and mIoU. These results highlight the strong generalization ability of AMD when facing novel datasets and manipulation tasks.
Generalization Assessment across MLLMs. To assess AMD’s generalization to different MLLM-generated narratives, we use four popular MLLMs, including Qwen-VL [2], X-InstructBLIP [29], LLaVA [20], and mPLUG-Owl [50], to regenerate test set narratives. AMD, trained on NYT, is then evaluated on the new test set, recording the TF category AP for both intra-domain (NYT) and cross-domain settings. As shown in Tab. 4b, bar chart I, AMD achieves intra-domain and cross-domain APs of over 76 and 53, respectively, demonstrating its robustness across diverse MLLMs.
@@ 5.2 Ablation Study
Tab. 4a presents the results for each component considered in our study. We use a fine-tuned Florence-2 with our designed prompts as the baseline for comparison. As shown, incorporating Artifact Pre-perception Encoding (APE) improves all three task metrics, especially binary classification accuracy, which increases from 76.92 to 82.93 on NYT and from 82.33 to 87.68 on Guardian. This demonstrates that pre-perception of manipulation traces is vital for aiding MLLMs in multi-media manipulation detection. Adding auxiliary tasks, such as Dual-Branch Manipulation (DBM) and Image Manipulation Grounding (IMG), enhances fake news classification and grounding performance, while also slightly improving binary classification. Notably, DBM significantly boosts AMD’s mAP, increasing from 47.18 to 66.47 on NYT and from 56.43 to 60.26 on Guardian. Furthermore, the incorporation of the Token Redundancy Penalty (TRP) yields comprehensive performance gains, especially exhibiting stable improvements in ACC and mIoU across both domains.

### v2

@@ 3.3 Dataset Statistics
With the above steps, we finally harvest our MDSM dataset $\mathcal{D}_{M}$, a large-scale, 100% semantic-aligned multi-modal benchmark with high-fidelity texts from MLLM. The distribution of manipulation categories is well balanced and consistent with previous datasets, ensuring fair evaluation across manipulation modes (Fig. 2a). Compared with the existing manipulation detection benchmarks in Tab. 1, MDSM has the following advantages: 1)Risk Consideration of MLLM. MDSM acknowledges the emerging challenges posed by MLLMs and utilizes multi-modal methods to create semantically coherent and contextually plausible narratives for manipulated images. This scenario, though underexplored, is a highly significant and timely problem in the modern large model era. 2)Semantic Alignment. MDSM is an aligned multi-modal media manipulation benchmark, which is a significant and more practical scenario for multi-modal manipulation detection. 3)Large Scale. Our MDSM comprises 441,423 samples and is the largest benchmark for detecting and grounding multi-modal manipulation. 4)Diverse Multi-media Sources. The multi-modal media of MDSM sources from diverse media sites, including The Guardian, The New York Times, The Washington Post, USA Today, and the BBC. Consequently, the generality of methods can be assessed via cross-domain evaluation.
@@ 4 Methodology
Figure 3: Overview of the proposed AMD framework. (a) Maps the manipulated image and prompts into a unified representation, incorporating an artifact token. (b) Utilizes the artifact pre-perception multimodal encoder $\mathcal{E}_{m}^{p}$ to extract perceptual clues. (c) Processes multi-modal features through $\mathcal{E}_{m}$ to generate text-based detection results. (d) Outputs and visualizes the final manipulation analysis.
Fig. 3 illustrates the architecture of our Artifact-aware Manipulation Diagnosis framework (AMD). Built upon Florence-2  [Florence2_CVPR] to leverage real-world knowledge, AMD follows a sequence-to-sequence framework for joint textual detection and grounding. Multimodal inputs are processed through three stages, and outputs localized predictions with textual explanations.
@@
In this prompt paradigm [^1] Details and examples are given in supplementary materials.,
  • $\textit{Task}>$: Specifies the manipulation detection objective and pairs the input image-text.
  • $\textit{Options}>$: Lists all candidate answers for MDSM task.
  • $\textit{Grounding}>$: Conditionally triggers region localization (via coordinates in brackets) only when image manipulation is detected.
  • $\textit{Response}>$: Encapsulates the correct answers.
  • $\textit{Coordinates}>]$: Optionally encloses tampered region coordinates if the image is altered.
Artifact Token Embeddings. To effectively adapt the MLLM into MDSM context while preserving its pretrained knowledge, we introduce a learnable Artifact Token that explicitly encodes artifacts from heterogeneous inputs. Formally, let the artifact token embeddings be denoted as ${E}_{a}\in\mathbb{R}^{n_{a}\times d}$, where $n_{a}$ indicates the token count and $d$ the embedding dimension. The textual input is processed through an embedding layer to obtain text embeddings ${E}_{t}\in\mathbb{R}^{n_{t}\times d}$ , while the visual input is encoded via a vision backbone $\mathcal{E}_{v}$ followed by a LayerNorm-augmented linear projection, yielding image embeddings ${E}_{v}\in\mathbb{R}^{n_{v}\times d}$. The above embeddings are concatenated to construct the input sequence: ${S}_{\text{inp}}=[{E}_{v};{E}_{a};{E}_{t}]$, where $\cdot$$\cdot$] means concatenating along the token dimension.
@@ 4.2 Artifact Pre-perception Encoding
This stage aims to perceive manipulation artifacts within input data and condense these forensic clues into the artifact token. Specifically, the input sequence $S$ undergoes processing through the artifact pre-perception multimodal encoder $\mathcal{E}^{p}_{m}$, yielding $\hat{S}=[\hat{E}_{v};\hat{E}_{a};\hat{E}_{t}]$. To inject artifact-aware clues into the artifact token embedding $\hat{E}_{a}$, we pick $\hat{E}_{a}$ from $\hat{S}$ and feed it into an artifact-aware classification head. As illustrated in Fig. 3, this classification head is optimized via a manipulation detection objective to explicitly encode artifact-related patterns into $\hat{E}_{a}$.
@@
$\mathcal{L}_{APE}=\mathbb{E}_{(I,T)\sim\mathcal{D}_{M}}\mathbf{CE}(C_{a}(\bar{E}_{a}),y_{fd}),$ (2)
@@ 4.3 Manipulation-Oriented Reasoning
Manipulation-Oriented Reasoning (MOR) is in charge of generating the textual answer in response to the question prompt. To acquire an accurate response, we augment the network optimization in MOR with two guiding tasks: visual Artifact Capture via Grounding and Manipulation-focused Guidance.
@@
Subsequently, the $u_{agg}$ is sent to the bbox detector to generate artifact coordinates. We follow [Loss_IoU] to construct the image manipulation grounding loss using L1 loss $\mathcal{L}_{1}$ and GIoU loss $\mathcal{L}_{IoU}$:
@@
Language modeling. The input sequence $S_{a}$ is processed through an encoder-decoder architecture, ultimately generating a pure text output that includes choices and coordinates (Fig. 3d) as specified in the prompts. In this stage, an autoregressive approach is adopted, where the decoder generates the target sequence $y$ conditioned on $S_{m}$. The language modeling loss $\mathcal{L}_{LM}$ [Florence2_CVPR] is used to supervise the decoded text outputs.
@@ 4.5 Training and Inference
5 Experiment
Table 2: Comparison of multi-modal learning methods on MDSM, where the background gray indicates the intra-domain performance. The better results in each group are in bold. AVG refers to the average performance across five news domains.
@@ 5 Experiment
Zero-Shot Qwen2.5-VL  [bai2025qwen25] 47.74 29.24 0.00 35.18 25.70 0.00 24.66 40.60 0.00 25.11 40.29 0.28 35.89 31.51 0.00 33.72 33.47 0.06
Qwen3-a22b  [Qwen3_blog] 45.29 25.01 0.71 38.12 27.41 1.19 22.87 39.17 0.10 22.87 40.77 1.34 37.22 31.60 0.98 33.27 33.69 0.86
GPT-4o  [hurst2024gpt4o] 48.48 27.90 0.82 35.68 29.49 1.23 24.62 39.88 1.37 23.62 38.89 1.22 37.19 30.48 1.20 33.92 33.33 1.17
Gemini-2.0  [team2023gemini] 56.05 33.16 1.44 41.26 24.37 1.12 29.60 38.29 1.40 29.15 35.20 2.42 38.12 29.13 2.25 38.83 32.03 1.72
Tr. on NYT ViLT  [ViLT_ICML] 83.27 64.27 22.73 72.18 31.76 20.21 70.34 36.45 21.48 65.71 36.23 17.56 74.33 36.10 19.36 73.17 40.96 20.27
HAMMER  [DGM4] 79.20 55.86 51.34 68.23 40.10 21.56 71.52 41.17 13.74 68.50 41.47 13.92 67.37 42.23 16.12 70.96 44.16 23.34
HAMMER++  [DGM4_TPAMI] 79.61 57.06 54.44 66.99 38.07 17.34 67.18 37.58 10.76 66.28 37.97 10.88 66.12 37.82 13.68 69.23 41.70 21.42
FKA-Owl  [liu2024fkaowl] 94.67 78.18 55.81 77.20 46.88 43.67 78.00 44.45 50.73 75.49 50.83 43.53 84.65 60.73 43.28 81.60 56.77 46.23
@@
Tr. on Guardian ViLT  [ViLT_ICML] 68.80 43.99 21.77 85.29 67.34 41.80 70.34 46.24 37.68 78.61 47.17 38.13 80.00 44.79 38.97 76.61 49.90 35.67
HAMMER  [DGM4] 61.89 37.98 18.84 78.50 52.40 51.53 74.78 50.76 43.40 75.11 50.34 46.36 81.32 50.15 56.03 74.32 48.33 43.23
HAMMER++  [DGM4_TPAMI] 62.75 36.45 23.76 80.95 59.92 64.67 75.36 48.77 47.13 76.30 49.56 48.91 80.12 50.36 57.97 75.10 49.01 48.49
FKA-Owl  [liu2024fkaowl] 80.60 40.44 26.33 92.60 78.24 71.04 80.90 51.80 50.93 78.88 51.62 50.88 87.61 68.57 61.82 84.12 58.13 52.20
@@
Table 3: Comparison of multi-modal learning methods on $\text{DGM}^{4}$, where the guardian domain with background gray is intra-domain. $\text{P}_{tok}$ is Precision of fake token grounding.
@@
ACC mAP $\text{P}_{tok}$ mIoU ACC mAP $\text{P}_{tok}$ mIoU ACC mAP $\text{P}_{tok}$ mIoU ACC mAP $\text{P}_{tok}$ mIoU ACC mAP $\text{P}_{tok}$ mIoU
ViLT  [ViLT_ICML] 68.27 42.29 69.87 43.19 52.79 31.28 62.11 33.78 55.76 33.26 57.17 31.10 44.14 39.68 59.06 21.96 55.24 36.63 62.05 32.49
HAMMER  [DGM4] 78.34 66.79 78.27 61.09 64.97 40.49 73.76 40.51 63.54 40.26 76.13 38.53 54.97 40.84 81.48 43.74 65.45 47.10 77.41 45.97
HAMMER++  [DGM4_TPAMI] 79.13 67.11 78.24 62.15 65.25 40.74 73.24 41.14 63.83 40.34 76.17 38.21 54.24 41.25 81.73 43.23 65.61 47.36 77.34 46.19
FKA-Owl  [liu2024fkaowl] 82.97 53.86 87.70 65.69 67.57 38.97 79.44 32.57 67.05 37.70 81.55 31.86 70.26 40.20 84.54 46.48 71.96 42.68 83.31 44.15
@@
Table 4: Ablation on (a) each components and (b) discussion regarding performance on test set of difference MLLMs & differnece manipulation type.
@@
Please refer to the appendix for the experimental setup and evaluation metrics.
5.1 Quantitative Results
Effectiveness & Generalization. We assess AMD against four SOTA methods on the MDSM and $\text{DGM}^{4}$ datasets. For MDSM (Tab. 2), we train on The Guardian and NYT, testing on the rest. For $\text{DGM}^{4}$ (Tab. 3), we train on the largest subset, The Guardian. Tab. 2 also shows zero-shot results for general-purpose models. Our key findings are: (1) MLLMs’ knowledge boosts performance. Forgery-trace methods like ViLT [ViLT_ICML] and HAMMER series [DGM4_TPAMI] show limited performance, unlike MLLM-based methods like FKA-Owl [liu2024fkaowl] and AMD. For instance, trained on MDSM-NYT (Tab. 2), AMD achieves an 84.47 intra-domain mAP and >60 cross-domain, while HAMMER scores 57.06 and <42, respectively. (2) AMD achieves strong grounding. AMD attains the best average mIoU of 63.56 (NYT-trained) and 61.02 (Guardian-trained) (Tab. 2). General-purpose models perform poorly (mIoU < 3). AMD’s superiority stems from its question-answer heuristic prompts and MOR module, which omits coordinate outputs when no manipulation is detected, thus reducing unnecessary errors. (3) AMD generalizes effectively. On $\text{DGM}^{4}$ (Tab. 3), AMD outperforms the HAMMER series on all metrics (74.47 ACC, 52.91 mAP, 80.01 $\text{P}_{tok}$, 51.87 mIoU). It also surpasses FKA-Owl in ACC, mAP, and mIoU, despite a lower $\text{P}_{tok}$.
Generalization Assessment across MLLMs. To assess generalization on different MLLMs, we evaluated an NYT-trained AMD on test narratives generated by four MLLMs: Qwen-VL [QwenVL], X-InstructBLIP [xInstructBLIP_2024], LLaVA [LLAVA_NeurIPS], and mPLUG-Owl [ye2024mplugowl3longimagesequenceunderstanding]. Results (Tab. 4b, chart I) show robust performance, with intra-domain (NYT) and cross-domain APs exceeding 76 and 53, respectively.
@@ 5.2 Ablation Study
Figure 4: Visualization of Artifact Token sequence. (a) Visualization of $E_{a}$ in $S_{inp}$. (b) Visualization of $\hat{E_{a}}$ in $S_{a}$. (c) Visualization of $\hat{E_{a}^{m}}$ in $S_{m}$
Component Ablation. Tab. 4a presents the results for each component considered in our study. We use a fine-tuned Florence-2 with our designed prompts as the baseline. As shown, incorporating Artifact Pre-perception Encoding (APE) improves all three task metrics, especially binary classification accuracy, which increases from 76.92 to 82.93 on NYT and from 82.33 to 87.68 on Guardian. This demonstrates that pre-perception of manipulation traces is vital for aiding MLLMs in multi-media manipulation detection. Adding auxiliary tasks, such as Dual-Branch Manipulation (DBM) and Image Manipulation Grounding (IMG), enhances fake news classification and grounding performance, while also slightly improving binary classification. Notably, DBM significantly boosts AMD’s mAP, increasing from 47.18 to 66.47 on NYT and from 56.43 to 60.26 on Guardian. Furthermore, the incorporation of the Token Redundancy Penalty (TRP) yields comprehensive performance gains, especially exhibiting stable improvements in ACC and mIoU across both domains.
Artifact Token Visualization. Fig. 4 visualizes the Artifact token at different stages via t-SNE [tsne_JMLR]. As shown in Fig. 4a to c, the sample points progressively form more distinct clusters, clearly demonstrating the effectiveness of our AMD optimization in enhancing the Artifact Token’s ability to distinguish between different categories.
For a more complete understanding of the MDSM dataset and the ablations on the AMD design, we refer readers to the appendix.